Back to the NIKA2 run 10 main page

April, 25th (Monday)

[Laurence 07h00-10h30]

At 6:30, the sky was found heavily cloudy and tau rising above 0.6. I restarted the acquisition and got ready waiting for any improvements of the weather. Atmospheric conditions were very unstable (tau225 jumping from 0.5 to 1) during the best part of the session. Around 9:50, the haze lifted and tau225 fell down quickly to about 0.3, then stabilized.

Mini Beam map session: The remaining 30 minutes were used to do a beam map on Neptune at the optimal focus (scan ID: 20170425s52)

Status of the beam map analysis can be seen at the bottom of the day-to-day log book page:

[Note to Olivier B.] the NIKA2 crate number 3 is powered using NIKA PWR 3. Juan had replaced the number-3 crate power by a spare (PWR 3) on Thursday at the beginning of the campaign.

Juan, Chris, Andrew & Hervé went down to Granada around 10:15 a.m. while Alessia attended a training on how to tune the pulse tube with Juan-Luis and with the help of Martino who has connected remotely.

Alessia & Laurence went down to Granada in the afternoon [TBC]

April, 24th (Monday)

[Chris & Laurence 08h00-16h00]

The tau has jumped to about 0.4 at the beginning of the shift, but then stabilized between 0.35 and 0.4. We had have to relaunch the acquisition after it crashed near 8:30. Since then, it worked without a glitch

We did a Beam map session: First we did optimal focus beammaps (scan ID: 20170424s116, 123, 134), as well as a defocus beammap sequence on Neptune (scan ID: 20170424s123-127). We did another optimal-z beammap on Mars (scan ID: 20170424s142) before the opacity rose and atmospheric day-time instabilities begun.

Preliminary analysis of the first beam map on Neptune at the optimal focus yields reasonable results

[Hervé & Juan 00h00-08h00]

At the beginning of our shift the weather improved significantly, the tau went down to 0.2 and stay between 0.25 and 0.3 for most of the night. We continue the cluster observations until 1h00 UTC. Then as the weather was good we decided to try Jean Francois's program after checking with him.

Here’s a preliminary reduction of the data:


Some results in M99 using all available scans:

m99_1mm_bis.jpeg m99_2mm.jpeg m99_spectralindex.jpeg

April, 23nd (Sunday)

[Andrew & Alessia 16h00-24h00]

As we took over the observations, the weather was extremely poor, with the taumeter reading 0.9 and extremely unstable. Poor conditions continued with us making occasional pointing and focus attemps to no avail, plus skydips until around 22h30 when the conditions improved dramatically. We quickly managed to get a good pointing and focus on the source 0716+714 close to the galaxy cluster, and we could then move on to making OTF maps for Science Verification.

[Chris & Laurence 08h00-16h00]

Observations started around 6:45 UTC, after the hail storm stopped.

We started doing a beam map session on Neptune. The off-line analysis of the first beam map scan at the optimal focus yields poor quality geomatries. Most of the kids of the 1mm arrays have to be discarded, which makes this defocused sequence of low utility..We moved to Mars at 8:30 UTC to get higher signal, even though at this time Mars is still at low-elevation (20-ish degree). We struggled for one hour and a half to get correct pointing and focus before stopping to observe. We checked from time to time for any improvements.

[Hervé & Juan 00h00-08h00]

Poor conditions, tau_225 around 0.5.

We did:

April, 22nd (Saturday)

[Andrew & Alessia 16h00-24h00]

The weather is terrible as we take over, with tau highly unstable around ~0.5-0.6. We tried a small pointing session to test the new Nasmyth offsets, but we struggled with all but the brightest sources available. We managed to get a good focus on 3C84, and pointing seemed to be stable for a little while. We moved over to take some observations of the SZ cluster, but the stability quickly disappeared and the weather worsened. Quite a while later, and after a few skydips, we resumed observations and tracked the two quasars 3C273 and 3C279 with a lack of any other options.

[Chris, Samuel & Laurence 08h00-16h00]

As soon as the wind felt down near 8 a.m., we started observing.

We did a complete beam map session: First a defocused beam map sequence on Neptune (scan ID 20170422s22, 23, 24, 31, 32). To follow the z-drift, the z-focus was estimated i) before the beam map sequence, ii) after 3 beam map scans and iii) after the sequence. The z drift was found to be rather modest, from -1.26 mm at the beginning of the sequance to -0.9 at the end. However, first analysis of the 2 optimal-focus beam maps (we ended the sequence with a second beam map at the optimal focus) indicates geometries of poor quality, with > 30% of the kids discarded in array 1. Then, we moved to Mars to have higher signal, and completed another defocused beam map sequence. Hovewer, we found a large z-focus drift of -0.9 mm between the beginning and the end of the sequence. We finished with a optimal-focus beam map at high elevation in poor weather conditions. The main goal was to test the PAKO script: the raw data file was correctly written on disk.

Around 1 p.m. UTC, Juan P. implemented new zero offset corrections into the pointing model. These corrections were independently derived by Albrecht, Nico and Juan P. We checked the pointing right after this change, and found small residual zero offsets to be set in PAKO (of 0.2, 0.2), which indicates the new model improves on the previous one.

[Hervé & Juan 00h00-08h00]

From the start of the session the wind speed is too high and we can not observe.



[Chris, Samuel & Laurence 08h00-16h00]

As soon as the wind felt down near 8 a.m., we started observing.

We did a complete beam map session: First a defocused beam map sequence on Neptune (scan ID 20170422s22, 23, 24, 31, 32). To follow the z-drift, the z-focus was estimated i) before the beam map sequence, ii) after 3 beam map scans and iii) after the sequence.

Then we did a gain elevation session on the secondary standards NGC7027, MWC349 and CRL2688 from 20 degree elevation to 75.

[Chris, Samuel & Laurence 00h00-08h00]

We continued a while the secondary calibrator session, adding a measurement at 80-ish degree of elevation. Howerver, we could do only one measure before the 3 calibrators reached the maximal elevation allowed for the telescope (of 83 degrees).

Then, we continued the pointing session, adding 7 points to improve a bit the coverage in azymuth. The results file has been updated and is available as /home/t22/Pointing_session_21_Apr_2017.odp

Then we went on to do a Beammap session. First we did a complete out-of-focus beammap sequence on Neptune. We note that the focus drifted from -0.75 at the beginning of the sequence to -0.25 at the end. This drift will make the exploitation of this defocused beammap series to estimate the focus surface tricky...Then, we did another beammap on Neptune at the optimal focus. We experienced difficulties to focus, likely to be due to poor and unstable weather conditions. The beammap scan (20170421s183) has to be cancelled after a crash of ELVIN. Alessia, Samuel and I work together to relaunch the acquisition and recover the resonances. A scan of Neptune beammap at optimal focus was redone, though in rather poor weather conditions.

[Alessia & Andrew 16h00-24h00]

Start of the observations with unstable atmospheric conditions. Tau=0.5-0.6. However after pointing and focus on a strong source we tried to observe the galaxy cluster but the wind went beyond the limit and we had to stop the telescope.

We did a preliminary analysis of the 20170419s133-137 defocus beammap sequence. Focus surface estimates are shown for the 3 arrays in the fig below. Each point is a focus measurement relative to the center using the closest kids in a radius of 40 arcsec. These results are consistent with focus surface estimates from NIKA2-Run9 data.


Albrecht has recomputed the pointing model from the pointing session. We do not update the pointing model.

April, 20th (Thursday)

[Hervé & Juan 00h00-08h00] We continue observing the cluster:

psz2g144_1mm.pdf psz2g144_2mm.pdf

Then, we started a session in M99 that lasted for about 2 hours of observation in total.

m99_1mm.pdf m99_2mm.pdf

We finally did another session of quasar observations to test the gain elevation curve.

[Samuel & Laurence, 8:00-16:00] Check list green (turbo pump: Tbottom about 25 deg Celsius; T mixing chamber about 182 milliK; cryo levels TBC; tunning OK)

Beammap session We did a defocused beammap sequence on Neptune beginning at an elevation of 30 deg: scan ID 20170420113-117. Then we did 4 beammaps at optimal focus on Mars at various elevation, namely 20170420s135 is around 30deg, 20170420s154 around 50 deg, 20170420s174 around 60 deg and 20170420s188 around 70 deg.

Preliminary gain vs elevation determination (NP): The first plot shows the flux of each source vs elevation for the currently available scans (no opacity correction applied but I assume a constant opacity and correct for the air mass). The second plot gathers all the sources vs elevation for each array. The fluxes are normalized to their value at 40 deg. elevation. The consistency so far is remarkable, thanks to constant X and Y focus throughout the observations ! On the second plot, the black line is a fit on all observations, the pink curve is the EMIR curve displayed for comparison. Although the fits look good, it's possible that further analysis (involving opacity correction) discards some of these scans, so if possible, it would be good to have one more sequence of monitoring these sources on the full range of elevations.

all_gain_vs_elevation_n2r10_v1.png all_gain_vs_elevation_n2r10_1_v1.png

[Alessia & Andrew, 16:00-00:00] As we start observing, everything on the instrumental side is working okay. As was the case yesterday, weather conditions are somewhat unstable at the start of the shift, despite blue skies outside, and tau is hovering around 0.36. We eventually managed to get focused and spend 1 hour doing a sequence on the galaxy cluster. After sunset, the weather worsened quite a bit, so science observations were out of the question. We tracked the two quasars 3C273 and 3C279 for some more OTF maps at different elevations.

April, 19th (Wednesday)

HA and JFMP [00h00:08h00]: During the night the average tau was 0.3 and very stable. We continue observing the two QSOs 3C273 and 3C279. We also did a beammap in 3C279. After that we switched to secondary calibrators (NGC7027, CRL2688 and MWC349) observations at different elevations. We repeated the same commands than for the QSO, doing OTF maps in azel and RADEC , and in az and elevation. We also performed some skydips.

[Samuel & Laurence, 8:00-16:00] check list: turbo pump working fine, mixing chamber temperature stable around 182 milliK, cryostat inclination stable, number of tuned kids good

Observations: We did a first beammap on Neptune (20170419s124), then 2 otf maps of Neptune for calibration stability checking while the temperature of the telescope stabilizes after the Sun has risen. We did a defocus beammap series on Neptune (20170419s133-137). Then we moved to Mars for having higher signal and did another beammap (20170419s151). The atmosphere begun to being very unstable at 11:30 UTC. We could not even focus on Mars. We moved to a small dark test session to check the f_tone values and then did a skydip.

Analyses: starting the analysis of 20170419s124. The antenna imbfits was missing, we generated it using the procedure described in The produced imbfits file generates error messages while producing the kidpar (nk_error: wrong imbfits file). We called Albrecht, who told us where to find the files on mrt-lx3 and scp-ed them to nika2-a imbfits_dir by hand. Later on, we learnt that launching nk_rta on each beammap scan would have done it...

Below you will find preliminary results on a geometry from a beammap scan on Neptune (scan ID: 20170419s133). This geometry (red ellipes) is compared with a geometry from one of our best beammap acquired during the 2 previous runs N2R8&9 (blue ellipses).

plot_compare_kidpars_20170419s133_A1.png plot_compare_kidpars_20170419s133_A3.png plot_compare_kidpars_20170419s133_A2.png

Average geometrical FWHM (from an 2D Gaussian fit) are:

kidpar: kidpar_20170419s133_v0

Note that beams are smaller than the N2R8 beams in the outter regions of the arrays (no beam broadening effect) whereas equivalent to N2R8 beams in the central region. This indicates better beams in average as expected when the focus is set to the best average value, i.e. best central value -0.2 mm.

[Alessia & Andrew, 16:00-00:00] Poor and unstable weather conditions. Tau=0.4. 18h : since it was not possible to observe, not possible to focus on Mars, Santiago went up in the cabin to open the plastic window on the vertex trying to refresh the cabin air a bit more. 20h : we finally found the best focus and we've started the observation of the galaxy cluster with an opacity almost stable to 0.3.

April, 18th (Tuesday)

Start of the official run.

First things first: verification of the alignment (since it changed for an unknown reason between run 8 and 9!)

==> ~2.5mm too high and ~1.5mm too much right w.r.t radio axis. Given the measurement errors and the expected effect on pupil illumination of the primary we think it's preferable to let it as it is. Indeed below you can see the Zemax simulations of illumination of the primary by various field taken over the FOV (various positions on the array) with a good alignment and with the offsets we measured. The effects on the beam are negligible (changes of structure at 10-3 level), it's also negligible for pointing. Zemax can't simulate the effects on the real data quality of the asymmetric illumination of the primary; we can just infer that as long as we stay inside M1 we should be fine, though we might have a small increase of background due to slightly higher spillover.

M1_all_beams_illum_without_mis-align_on_Widow.PNG M1_central_beam_illum_with_2p5mmUp_1p5mmRight_mis-align_on_Widow.PNG

AlR,AR [16h00-00h00]: We start the observations in poor weather (tau=0.5), and without a properly functioning turbo pump. Since Mars is bright, we intended to follow it for some beam maps, though by the time we managed to get focused, the turbo pump was having trouble and the cabin was too warm again. Once the cabin was cool again, Mars was too low, so we moved onto the quasar 3C273, though it was too faint at 1mm to make any sensible conclusions about the focus. After this, we targeted the nearby quasar 3C279, which is a bit bright, and were able to get focused successfully. Given the weather conditions and the lack of other bright targets, we carried out OTF maps of the QSOs 3C273 and 3C279 at different elevations.

April, 14th (Friday)

Alessia at the telescope and Juan remotely.

We started doing two dark scans. We tried to point using 3C84 but the source was far from the center -38,-20 arcsec. After discussions with Juan Peñalver we decided to use nasmyth offset computed by Nicolas on run 9: offsets -18.8 -5 /sys nasmyth, However, this did not solve either the problem. We switched to Mars but encountered the same problem. Finally, we found that the problem was related to a wrong reference kidpar - we corrected this and everything started work fine. However the weather during the day was very unstable although tau was not that bad: about 0.3-0.4 from IRAM taumeter. We have performed a small pointing session to verify the current model. We tried to get some Mars beammaps after focus X,Y, Z. We did a couple of skydips. Finally, we tried to go to PSZ2G144.85 to do a first test but had no enough time. Really a pity because the weather improved quite a lot.

Alessandro 22h45. I stop the turbo for the weekend. Plan to switch it ON on Monday in the late evening. Alessia, could you put the NIKA2 window back in place so we can check the noise between Monday and Tuesday ? Good night !

DailyReportsNika2Run10 (last edited 2017-04-25 09:43:34 by NikaBolometer)